For the first 50 miles, Hannah traveled at 25 miles per hour. Then she doubled her speed. If the entiretrip was 450 miles, how long did the trip take?
Answer by stanbon(75887)

For the first 50 miles, Hannah traveled at 25 miles per hour. Then she doubled her speed. If the entiretrip was 450 miles, how long did the trip take?
-------
1st leg DATA:
dist = 50 miles ; rate = 25 mph ; time = d/r = 50/25 = 2 hrs
-------------------
2nd leg = 400 miles ; rate = 50 mph ; time = d/r = 400/50 = 8 hrs
-----
Total time = 2hr + 8hr = 10 hours
-----
Cheers,
Stan H.
-------------